OpenStudy (anonymous):

SOMEONE HELP PLEASEEEE Suppose the volume of a given object varies directly as the cube of the radius of its base. If one example of the object with radius 4.5 feet has a volume of 127.2 ft3, what is the volume of a second example with radius 2.5 feet?

OpenStudy (anonymous):

use proportions

OpenStudy (anonymous):

I think it's 57.24

OpenStudy (anonymous):

A. 70.67 feet3 B. 42.27 feet3 C. 39.25 feet3 D. 21.81 feet3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

>.< mhm im stuck

OpenStudy (phi):

volume varies directly as the cube of the radius that means V= k*r^3 (k is a number) they give us V and r^3 for one of the objects, so we can find k 127.2= k*(4.5)^3 can you find k?

OpenStudy (anonymous):

no lol im stuck. >.<

OpenStudy (phi):

do you have a calculator? can you find 4.5^3 (4.5 cubed, which means 4.5*4.5*4.5)

OpenStudy (anonymous):

91.125

OpenStudy (phi):

127.2= k*(4.5)^3 can you rewrite this equation, replacing 4.5^3 with its value (the number you found)

OpenStudy (anonymous):

127.2=k*(91.125) ?

OpenStudy (phi):

yes. now to "solve for k". divide both sides by the number multiplying k: 127.2/91.125= k*91.125/91.125 on the right side, any number divided by itself is 1. Can you rewrite the equation (do the division on both sides)

OpenStudy (anonymous):

1.39 and a whole bunch more number ?

OpenStudy (phi):

yes, but write it as an equation. start with 127.2/91.125= k*91.125/91.125 what do we get ?

OpenStudy (anonymous):

1.39=k ?

OpenStudy (phi):

yes. so k = 1.39 (roughly, because we dropped some numbers) but let's see what we get when we use it. V= k*r^3 (volume varies directly as the radius cubed is how you interpret this) using k = 1.39 we have V= 1.33*r^3 second example with radius 2.5 feet, replace r in the equation with the number 2.5 what do we get? write down the equation

OpenStudy (phi):

*1.39 not 1.33

OpenStudy (anonymous):

V=21.71

OpenStudy (phi):

which should be close to one of your choices. If you recompute k with say 4 decimals, the answer will match

OpenStudy (anonymous):

21.81?

OpenStudy (phi):

yes. if you used k= 1.3959 (rounded) you would get 2.81 (rounded to the nearest 100th)

OpenStudy (anonymous):

okay thanks hun <3

OpenStudy (phi):

if you want more info on how to do the algebra, try khan's site: http://www.khanacademy.org/math/algebra/solving-linear-equations/v/simple-equations it is not too hard to learn this ...